OpenStudy (anonymous):

Simplify. Identify any x-values for which the expression is undefined. \[\frac{ 6x ^{2}+7x-3 }{ -3x ^{2}+x }\]

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

ok first x values for which the expression is undefined in order to find them we have to find the x values who makes the denominator 0 -3x^2 +x = 0 can you find them ?

OpenStudy (anonymous):

Im not very good at this and am stuck thats the only reason im on here cause i have no clue and school in two hours

OpenStudy (anonymous):

ok so in order to solve -3x^2 + x =0 we take out one x since in 2 of the terms there is x and get : x (-3x + 1) = 0 now two solutions for this x=0 or -3x +1 = 0 and it means x = 1/3 so for x = 0 and x = 1/3 the expression is undifined

OpenStudy (anonymous):

now the numerator can be simplify into 6x^2 +7x - 3 6x^2 +9x-2x - 3 6x(x+1.5)-2(x+1.5) (6x-2)(x+1.5) 2(3x-1)(x+1.5) so we have : [2(3x-1)(x+1.5)] / [x(-3x + 1)] and you should know that (3x-1) / (-3x+1) = -1 so : -2(x+1.5) / x this is the simplified form

OpenStudy (anonymous):

thank you so much!! you showed just enough work for me to understand how you did that that was great thank you

OpenStudy (anonymous):

yw one more thing to say here is 6x^2 +7x - 3 do you know how i chose to split 7x into 9x-2x ?

OpenStudy (anonymous):

so you could split it into the ( )

OpenStudy (anonymous):

yes but how i knew to split it exacly to 9x-2x and not 5x+2x for example if you know why it's good if not ill explain